OpenStudy (anonymous):

use quadratic formula .3x^2=.6x+.3 use integers or fractions

OpenStudy (anonymous):

as whole numbers or fractions?

OpenStudy (anonymous):

-.3x^2 + .6x + .3 = 0 a= -.3 b= .6 c= .3 Use quadratic formula to solve for x. \[-b \pm \sqrt{b^2-4ac?}/2a\]

OpenStudy (anonymous):

Ignore the ? in the formula, don't know why that showed up. The equation was correct when I hit insert.

OpenStudy (anonymous):

Your correct answers should be x= -0.0373, -0.7243. Hope this helped!

OpenStudy (anonymous):

how do u transfer that into a fraction

OpenStudy (anonymous):

-.0373 would just be -373/10000 and -0.7243 would be -7243/10000. They don't reduce.

OpenStudy (anonymous):

what if we kept it in radical terms